Reports 1-1 of 1 Clear search Modify search
AdV-PSL (Laser Bench setup)
derossi, spinicelli, chiummo We - 11:56 Friday 14 May 2021 (51786) Print this report
Comment to swap ML (51750)

**** tuning of the ML crystal temperature ****

In order to correct the bimodal behaviour of the ML (1st plot. We were at 24.76 C and at 25.55 C we started to observe the second mode) we tuned the temperature of the ML crystal. Following the procedure (50944) we found mode hopes appearing between 20.88C- 21.4 C,  24.1 C - 24.55 C and 27.3 C- 27.85 C.

We lowered the temperature from 24.76 C to 23 C (2nd plot).


**** increasing of the fiber amplfier seeder power ****

Following Walid's advice we increased the seeder power from 40 mW to 70 mW, which corresponds to 8V on the photodiode (3rd plot), in order to have more margin for seeding and avoid switch off if any polarization  fluctuations or small misalignment occur.

We switched off the fiber amplifier and turned the waveplate before teh pickoff cube from 70 deg to 77 deg. The SL still locks.

We also measured the power at the output of the ML fiber and we have 175 mW.

After having turned on again the fiber amplifier we relocked the injection.

Once all the loops closed an unlock of the PMC occurred but we didn't identify its cause. It relocked by itself few minutes later and until now it seems to be stable. We keep monitoring the SL behaviour because it could be that we ara close to the threshold for unlocking and whenever it unlocks some light could perturb the fiber amplifier or the ML. It this case a solution will be to keep the SL unlocked and put a beam dump at its output.

Images attached to this comment
Search Help